In a test of

Upper H 0H0​:

muμequals=100

against

Upper H Subscript aHa​:

muμnot equals≠​100,

the sample data yielded the test statistic

z equals 2.16z=2.16.

Find the

Upper PP​-value

for the test.
Answer #1

this is the two tailed test .  

The null and alternative hypothesis is ,

H0 :   = 100

Ha :     100

Test statistic = z=2.16

P(z >2.16 ) = 1 - P(z < 2.16 ) = 1 -0.9846=0.0154

P-value = 2*0.0154=0.0308